Find the standard form f(x)=a(x-h)^2+k of the quadratic function that has a y-intercept of 8 and the x-intercepts of 8 and -8, if the vertex and the y-inetercept are the same point.
f(x)=

Answer by htmentor(1343)

The y-intercept is the point (0,8), which is also the vertex.
So f(x) = a(x-0)^2 + 8
We know the function passes through the two points (-8,0) and (8,0). Use one of these to find a.
0 = a*8^2 + 8 -> a = -1/8
Thus f(x) = (-1/8)x^2 + 8